OpenStudy (cypress7):

I will Medal! what is 4 1/5 + 3 1/10

OpenStudy (anonymous):

you have : 41/5: Multiply and divide by 2, so that you get 10 in denominator: \[\frac{41}{5} \times \frac{2}{2} + \frac{31}{10} \implies \frac{82}{10} + \frac{31}{10}\]

OpenStudy (anonymous):

ok do u have a caculator?

OpenStudy (anonymous):

if u do the on ur caculator type in 4.20+3.10 ok and tell me wat u get .....

OpenStudy (anonymous):

As denominators are same now, so just add then numerator.. :)

OpenStudy (cypress7):

Okay thank you

OpenStudy (anonymous):

you welcome :) XD

OpenStudy (anonymous):

\[\frac{a}{c} + \frac{b}{c} = \frac{a+b}{c}\]